# TESLA-X: An effective method to search for sub-threshold lensed gravitational waves with a targeted population model

### v2

@@ IV.1 The missing pieces in the traditional TESLA method: Non-optimal template bank and population model
IV.2 The likelihood ratio in the GstLAL search pipeline
To rank candidates, GstLAL assigns each trigger a likelihood ratio as its ranking statistic, defined as Sachdev et al. (2019); Cannon et al. (2021); Messick et al. (2017):
@@ IV.2 The likelihood ratio in the GstLAL search pipeline
The method of calculating the numerator and denominator are described in detail in Fong (2018) and Cannon et al. (2015), respectively. While the derivations are not repeated here, we wish to highlight that information about the population model is enfolded into the numerator; therefore, it is $P(\vec{\theta}|\mathrm{signal})$ that allows the pipeline to perform targeted source searches. The following section describes the changes made to improve the strongly-lensed gravitational-wave search.
IV.3 A better reduced template bank, and a lensing-based targeted population model
Traditional TESLA aims to reduce the noise background effectively to uncover possible strongly-lensed counterparts to the target super-threshold gravitational wave. We reduce the number of templates, and hence trials factors, by keeping only templates that can cover the parameter space where possible lensed counterparts may live in. Through an injection campaign, traditional TESLA locates a subset of templates from the full template bank to construct a reduced template bank to search for possible lensed counterparts to a given target super-threshold gravitational wave. However, as explained at the beginning of IV, there are two problems in traditional TESLA, both of which lead to a decrease in the search sensitivity of TESLA.
@@ IV.3 A better reduced template bank, and a lensing-based targeted population model
Under the strong lensing hypothesis, possible lensed counterparts should have the same parameters as the target super-threshold gravitational waves regardless of being super-threshold or sub-threshold. Therefore, the population of possible lensed counterparts to the target gravitational waves should have a higher probability of lying near the signal subspace defined by the posterior parameter probability distribution of the super-threshold event, and a lower probability of living far away from that. When we perform the injection campaign in traditional TESLA, we can obtain both the recovered injections’ SNRs and the corresponding templates that recover the injections. Generally speaking, sub-threshold signals with lower optimal SNRs are affected by noise fluctuations more substantially and hence can be recovered by templates with parameters different from those of the target super-threshold event. That said, the templates closer to the signal subspace defined by the posterior parameter probability distribution of the target are associated with recovered injections with higher SNR values. Therefore, we can make use of the results from the injection campaign to construct a Gaussian Kernel Density Estimation (KDE) function $\text{KDE}(\vec{\gamma})$ using the distribution of the templates that recovered the lensed injections, with the corresponding injections’ recovered SNRs as weights. The KDE is a function of source parameters $\vec{\gamma}$, which, to match the parameters of the template bank, are set to be the masses and spins of the source. When weighted with the recovered SNRs, regions closer to the signal subspace that the target super-threshold gravitational wave lies in will give a higher value for the KDE function, which matches our expectation that lensed counterparts to the target should have similar parameters as the target. The KDE function can then be redefined as the probability density function that describes the “source population” for our possible strongly-lensed sub-threshold counterparts to the target gravitational wave. Using this as our population model, we then follow the steps detailed in Fong (2018) to calculate $P(\vec{\theta}|\mathrm{signal})$, i.e. the probability that a signal is recovered by a template with parameters $\vec{\theta}$. Notably, this probability depends on SNR, such that the search does not overly penalize templates far away from the signal subspace for weak sub-threshold signals. These methods instill a correct population model to the new TESLA method compared to the incorrect uniform population model in traditional TESLA, improving the search sensitivity.
IV.4 The flow of the TESLA-X search method
@@
After the injection campaign, we now retrieve a set of templates that can recover the simulated lensed injections, and the associated observed SNRs of the recovered injections. We can plot the retrieved templates on the chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ space, with the colors of the points being the associated recovered injections’ observed SNRs $\rho_{\text{obs}}$. The top left panel of Figure 2 shows an example cartoon that plots the templates on the $\mathcal{M}_{c}\mathrm{-}\chi_{\text{eff}}$ space, with the color being the recovered injections’ network SNRs.
@@
Because the Gaussian KDE function $f(\vec{\gamma})$ is essentially a probability density function (PDF) that describes the expected distribution of possible (sub-threshold) lensed counterparts to the target super-threshold gravitational wave, we can use it as the astrophysical source population prior to construct a targeted population model using the same methodology described in the previous section and Fong (2018). Figure 3 shows an example cartoon to illustrate the idea.
@@
V Method verification and results
In Li et al. (2023), we carry out a full-scale simulation campaign, where we prepare mock data with a pair of super-threshold and sub-threshold lensed signals injected, and apply the TESLA method to recover the two signals. We have proved that the TESLA method can recover sub-threshold lensed signals if they exist. TESLA-X and TESLA have essentially the same working principle, but the only differences are how we construct the reduced template bank and the population model. Therefore, we conduct a simulation campaign solely to compare the search effectiveness and sensitivity towards possible sub-threshold lensed gravitational-wave signals using TESLA and the proposed TESLA-X method.
V.1 Mock data generation
First, we generate a $9.15$-day long data stream with Gaussian noise recolored with O4 characteristic power spectral densities for LIGO Hanford, LIGO Livingston, and Virgo detector. The exact data for O4 characteristic power spectral densities for LIGO Hanford, LIGO Livingston and Virgo detector can be found at https://dcc.ligo.org/T2200043-v3/public. We assume that there is no detector downtime (Note: A detector is considered “down” if it is not in observing mode.), and no times are vetoed. We inject a super-threshold gravitational wave generated using the IMRPhenomXPHMpseudoFourPN waveform Pratten et al. (2021) into the mock data. Details about the source parameters of the injected gravitational wave are listed in Table 1.
@@ V.1 Mock data generation
V.2 Performing a general search
We then use GstLAL to perform a search over the mock data stream following the same settings used to search for gravitational waves within O3 data in GWTC-3 Abbott et al. (2021b). We use the same general template bank as described in Li et al. (2023) for the general search. As expected, the search recovers MS220508a with the highest ranking statistics (FAR$=2.972\times 10^{-35}$ Hz, rank 1) among all other triggers. We then apply a Bayesian inference library for gravitational-wave astronomy Bilby Ashton et al. (2019) 111 For reference, we used the default MCMC approach to do the sampling. to perform parameter estimation (PE) for MS220508a, which outputs a set of posterior samples required to apply the TESLA and TESLA-X search pipeline.
V.3 Performing an injection campaign
@@
In the injection campaign, the general template bank recovers $6151$ out of $7815$ injections (See the column “General” in Table 2), ringing up a total of $1008$ templates (Note: The same template may be rung up multiple times by different injections, and hence the number of rung-up templates can be fewer than the number of found injections.). Figure 4 plots the rung-up templates on the chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ space, with the colors of the markers being the network SNRs of the associated injections. We can see that the results match our expectations: Templates closer to the actual parameters of MS220508a are associated with injections with higher recovered SNRs (i.e. super-threshold). Those that have parameters much different than the actual parameters of MS220508a are, in general, related to injections with lower SNRs (i.e. sub-threshold).
Figure 4: Rung-up templates by the simulated lensed injections of MS220508a in the injection campaign, plotted on the chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ space. The colors of the markers represent the network SNRs of the associated injections.
V.4 Generating the TESLA and TESLA-X reduced template banks
@@
Figure 5: The templates in the full template bank (in grey) and in the TESLA reduced template bank (in orange), plotted in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space.
@@
Figure 6: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220508a in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space. The colors represent the probability density estimated by the KDE function.
We then take the lowest contour level (For the case of MS220508a, the lowest contour level has a value of $0.005$) as a boundary, and use all the templates within this boundary to construct a TESLA-X targeted template bank. Figure 7 shows the templates in the full template bank (in grey) and in the TESLA-X reduced template bank (in red) in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space. The yellow curve marks the lowest contour level we obtained for the estimated Gaussian KDE function we showed in Figure 6. There are $22136$ templates in the TESLA-X reduced template bank.
Figure 7: The templates in the full template bank (in grey) and in the TESLA-X reduced template bank (in red), plotted in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space. The TESLA-X template bank is constructed by keeping templates that fall within the lowest contour level from the Gaussian KDE function shown in Figure 6, i.e. $0.005$, plotted as a yellow curve in this figure. For easy comparison, we plot the TESLA reduced template bank templates in the same figure (in blue).
V.5 Constructing the targeted population model for the TESLA-X reduced template bank
Next, we use the estimated KDE function for MS220508a and the TESLA-X reduced template bank to construct a targeted population model, following the same steps detailed in IV and Fong (2018). Note that for practical reasons, the Gaussian KDE function $\text{KDE}(\vec{\gamma})$ is re-calculated with $\vec{\gamma}$ chosen to be the component masses and spins $m_{1},m_{2},\chi_{1},\chi_{2}$ instead of $\mathcal{M}_{c}$ and $\chi_{\text{eff}}$.
V.6 Performing a re-filtering to recover the lensed injections with the TESLA and TESLA-X methods
@@
Figure 8: (Top panel) The sensitive range v.s. FAR for MS220508a-alike signals using the full template bank (black), TESLA template bank (red), and TESLA-X template bank (blue), respectively. The shaded band for each curve represents the corresponding $1$-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that both the TESLA and TESLA-X bank improves in terms of sensitivity towards MS220508a-alike (lensed) signals, but the curve representing the TESLA-X bank (blue) is above that of the TESLA bank (red), meaning that the TESLA-X method outperforms the TESLA method. In particular, notice that the improvement is most significant in the region with FAR $\geq 10^{-7}\text{ Hz}$, which represents the “sub-threshold” region. This further proves that the TESLA-X method is better than the traditional TESLA method.
@@
VI Concluding Remarks
The detectability of strongly lensed gravitational wave (GW) signals can be compromised as they may experience demagnification, falling below the detection threshold. In a previous study, we introduced the TESLA method Li et al. (2023) to extract strongly lensed sub-threshold GW signals by constructing a target template bank based on the information from detected events. This paper presents an enhanced approach called TESLA-X, which aims to improve the search sensitivity for strongly lensed sub-threshold GW signals.
@@ VI Concluding Remarks
The targeted population model is constructed through several steps. Initially, a Gaussian kernel density estimation (KDE) function is applied to the retrieved injections’ source parameters $\vec{\gamma}$. This KDE function is then transformed into a probability density function, representing the parameter space of our potential strong-lensed sub-threshold GW signals Fong (2018). Additionally, we generate the targeted template bank by defining a boundary within the retrieved injections. This process involves discarding templates from the original full template bank outside of a predefined boundary that encloses the target parameter space while keeping all templates within the boundary. This results in a densely distributed template bank targeting a smaller region. Compared to the TESLA method, this refined template bank ensures the recovery of any lensed counterparts with less than a $3\%$ SNR loss. Notably, the FARs and likelihood values ($\mathcal{L}$) do not provide information about the likelihood of being lensed counterparts to the target event. Instead, they serve as metrics to differentiate between noise triggers and astrophysical signals.
@@
Acknowledgements.
The authors acknowledge the generous support from the National Science Foundation in the United States. AKYL would like to gratefully acknowledge the sup- port from the National Science Foundation through the Grants NSF PHY-1912594 and NSF PHY-2207758. JCLC acknowledges support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. HF gratefully acknowledges the support of the Canadian Institute for Theoretical Astrophysics (CITA) National Fellowship and the Japan Society for the Promotion of Science (JSPS) Postdoctoral Fellowships for Research in Japan. JME is supported by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 INTERACTIONS, and by VILLUM FONDEN (grant no. 53101 and 37766). The authors are also grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org) Vallisneri et al. (2015), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-0757058. HF is supported by a CITA National Fellowship. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes. This paper carries LIGO Document Number LIGO-P1900059.

### v5

@@ 4.1 The missing pieces in the traditional TESLA method: Non-optimal template bank and population model
4.2 The likelihood ratio in the GstLAL search pipeline
To rank candidates, GstLAL assigns each trigger a likelihood ratio as its ranking statistic, defined as (Sachdev et al., 2019; Cannon et al., 2021; Messick et al., 2017):
@@ 4.2 The likelihood ratio in the GstLAL search pipeline
The method of calculating the numerator and denominator are described in detail in (Fong, 2018) and (Cannon et al., 2015), respectively. While the derivations are not repeated here, we wish to highlight that information about the population model is enfolded into the numerator; therefore, it is $P(\vec{\theta}|\mathrm{signal})$ that allows the pipeline to perform targeted source searches. The following section describes the changes made to improve the strongly-lensed gravitational-wave search.
4.3 A better reduced template bank, and a lensing-based targeted population model
Traditional TESLA aims to reduce the noise background effectively to uncover possible strongly-lensed counterparts to the target super-threshold gravitational wave. We reduce the number of templates, and hence trials factors, by keeping only templates that can cover the parameter space where possible lensed counterparts may live in. Through an injection campaign, traditional TESLA locates a subset of templates from the full template bank to construct a reduced template bank to search for possible lensed counterparts to a given target super-threshold gravitational wave. However, as explained at the beginning of 4, there are two problems in traditional TESLA, both of which lead to a decrease in the search sensitivity of TESLA.
@@ 4.3 A better reduced template bank, and a lensing-based targeted population model
Under the strong lensing hypothesis, possible lensed counterparts should have the same parameters as the target super-threshold gravitational waves regardless of being super-threshold or sub-threshold. Therefore, the population of possible lensed counterparts to the target gravitational waves should have a higher probability of lying near the signal subspace defined by the posterior parameter probability distribution of the super-threshold event, and a lower probability of living far away from that. When we perform the injection campaign in traditional TESLA, we can obtain both the recovered injections’ SNRs and the corresponding templates that recover the injections. Generally speaking, sub-threshold signals with lower optimal SNRs are affected by noise fluctuations more substantially and hence can be recovered by templates with parameters different from those of the target super-threshold event. That said, the templates closer to the signal subspace defined by the posterior parameter probability distribution of the target are associated with recovered injections with higher SNR values. Therefore, we can make use of the results from the injection campaign to construct a Gaussian Kernel Density Estimation (KDE) function $\text{KDE}(\vec{\gamma})$ using the distribution of the templates that recovered the lensed injections, with the corresponding injections’ recovered SNRs as weights. The KDE is a function of source parameters $\vec{\gamma}$, which, to match the parameters of the template bank, are set to be the masses and spins of the source. When weighted with the recovered SNRs, regions closer to the signal subspace that the target super-threshold gravitational wave lies in will give a higher value for the KDE function, which matches our expectation that lensed counterparts to the target should have similar parameters as the target. The KDE function can then be redefined as the probability density function that describes the “source population” for our possible strongly-lensed sub-threshold counterparts to the target gravitational wave. Using this as our population model, we then follow the steps detailed in (Fong, 2018) to calculate $P(\vec{\theta}|\mathrm{signal})$, i.e. the probability that a signal is recovered by a template with parameters $\vec{\theta}$. Notably, this probability depends on SNR, such that the search does not overly penalize templates far away from the signal subspace for weak sub-threshold signals. These methods instill a correct population model to the new TESLA method compared to the incorrect uniform population model in traditional TESLA, improving the search sensitivity.
4.4 The flow of the TESLA-X search method
@@
After the injection campaign, we now retrieve a set of templates that can recover the simulated lensed injections, and the associated observed SNRs of the recovered injections. We can plot the retrieved templates on the chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ space, with the colors of the points being the associated recovered injections’ observed SNRs $\rho_{\text{obs}}$. The top left panel of Figure 2 shows an example cartoon that plots the templates on the $\mathcal{M}_{c}\mathrm{-}\chi_{\text{eff}}$ space, with the color being the recovered injections’ network SNRs.
@@
Because the Gaussian KDE function $f(\vec{\gamma})$ is essentially a probability density function (PDF) that describes the expected distribution of possible (sub-threshold) lensed counterparts to the target super-threshold gravitational wave, we can use it as the astrophysical source population prior to construct a targeted population model using the same methodology described in the previous section and (Fong, 2018). Figure 3 shows an example cartoon to illustrate the idea.
@@
5 Method verification and results
In (Li et al., 2023), we carry out a full-scale simulation campaign, where we prepare mock data with a pair of super-threshold and sub-threshold lensed signals injected, and apply the TESLA method to recover the two signals. We have proved that the TESLA method can recover sub-threshold lensed signals if they exist. TESLA-X and TESLA have essentially the same working principle, but the only differences are how we construct the reduced template bank and the population model. Therefore, we conduct a simulation campaign solely to compare the search effectiveness and sensitivity towards possible sub-threshold lensed gravitational-wave signals using TESLA and the proposed TESLA-X method.
5.1 Mock data generation
First, we generate a $9.15$-day long data stream with Gaussian noise recolored with O4 characteristic power spectral densities for LIGO Hanford, LIGO Livingston, and Virgo detector. The exact data for O4 characteristic power spectral densities for LIGO Hanford, LIGO Livingston and Virgo detector can be found at https://dcc.ligo.org/T2200043-v3/public. We assume that there is no detector downtime (Note: A detector is considered “down” if it is not in observing mode.), and no times are vetoed. We inject a super-threshold gravitational wave generated using the IMRPhenomXPHMpseudoFourPN waveform (Pratten et al., 2021) into the mock data. Details about the source parameters of the injected gravitational wave are listed in Table 1.
@@ 5.1 Mock data generation
5.2 Performing a general search
[…]
5.3 Performing an injection campaign
@@
In the injection campaign, the general template bank recovers $6151$ out of $7815$ injections (See the column “General” in Table 2), ringing up a total of $1008$ templates (Note: The same template may be rung up multiple times by different injections, and hence the number of rung-up templates can be fewer than the number of found injections.). Figure 4 plots the rung-up templates on the chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ space, with the colors of the markers being the network SNRs of the associated injections. We can see that the results match our expectations: Templates closer to the actual parameters of MS220508a are associated with injections with higher recovered SNRs (i.e. super-threshold). Those that have parameters much different than the actual parameters of MS220508a are, in general, related to injections with lower SNRs (i.e. sub-threshold).
Figure 4: Rung-up templates by the simulated lensed injections of MS220508a in the injection campaign, plotted on the chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ space. The colors of the markers represent the network SNRs of the associated injections.
5.4 Generating the TESLA and TESLA-X reduced template banks
@@
Figure 5: The templates in the full template bank (in grey) and in the TESLA reduced template bank (in orange), plotted in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space.
@@
Figure 6: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220508a in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space. The colors represent the probability density estimated by the KDE function.
We then take the lowest contour level (For the case of MS220508a, the lowest contour level has a value of $0.005$) as a boundary, and use all the templates within this boundary to construct a TESLA-X targeted template bank. Figure 7 shows the templates in the full template bank (in grey) and in the TESLA-X reduced template bank (in red) in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space. The yellow curve marks the lowest contour level we obtained for the estimated Gaussian KDE function we showed in Figure 6. There are $22136$ templates in the TESLA-X reduced template bank.
Figure 7: The templates in the full template bank (in grey) and in the TESLA-X reduced template bank (in red), plotted in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space. The TESLA-X template bank is constructed by keeping templates that fall within the lowest contour level from the Gaussian KDE function shown in Figure 6, i.e. $0.005$, plotted as a yellow curve in this figure. For easy comparison, we plot the TESLA reduced template bank templates in the same figure (in blue).
5.5 Constructing the targeted population model for the TESLA-X reduced template bank
Next, we use the estimated KDE function for MS220508a and the TESLA-X reduced template bank to construct a targeted population model, following the same steps detailed in 4 and (Fong, 2018). Note that for practical reasons, the Gaussian KDE function $\text{KDE}(\vec{\gamma})$ is re-calculated with $\vec{\gamma}$ chosen to be the component masses and spins $m_{1},m_{2},\chi_{1},\chi_{2}$ instead of $\mathcal{M}_{c}$ and $\chi_{\text{eff}}$.
5.6 Performing a re-filtering to recover the lensed injections with the TESLA and TESLA-X methods
@@
Figure 8: (Top panel) The sensitive range v.s. FAR for MS220508a-alike signals using the full template bank (black), TESLA template bank (red), and TESLA-X template bank (blue), respectively. The shaded band for each curve represents the corresponding 11-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that both the TESLA and TESLA-X bank improves in terms of sensitivity towards MS220508a-alike (lensed) signals, but the curve representing the TESLA-X bank (blue) is above that of the TESLA bank (red), meaning that the TESLA-X method outperforms the TESLA method. In particular, notice that the improvement is most significant in the region with FAR $\geq 10^{-7}\text{ Hz}$, which represents the “sub-threshold” region. This further proves that the TESLA-X method is better than the traditional TESLA method.
@@
6 Concluding Remarks
The detectability of strongly lensed gravitational wave (GW) signals can be compromised as they may experience demagnification, falling below the detection threshold. In a previous study, we introduced the TESLA method (Li et al., 2023) to extract strongly lensed sub-threshold GW signals by constructing a target template bank based on the information from detected events. This paper presents an enhanced approach called TESLA-X, which aims to improve the search sensitivity for strongly lensed sub-threshold GW signals.
@@ 6 Concluding Remarks
The targeted population model is constructed through several steps. Initially, a Gaussian kernel density estimation (KDE) function is applied to the retrieved injections’ source parameters $\vec{\gamma}$. This KDE function is then transformed into a probability density function, representing the parameter space of our potential strong-lensed sub-threshold GW signals (Fong, 2018). Additionally, we generate the targeted template bank by defining a boundary within the retrieved injections. This process involves discarding templates from the original full template bank outside of a predefined boundary that encloses the target parameter space while keeping all templates within the boundary. This results in a densely distributed template bank targeting a smaller region. Compared to the TESLA method, this refined template bank ensures the recovery of any lensed counterparts with less than a $3\%$ SNR loss. Notably, the FARs and likelihood values ($\mathcal{L}$) do not provide information about the likelihood of being lensed counterparts to the target event. Instead, they serve as metrics to differentiate between noise triggers and astrophysical signals.
@@
Acknowledgements
The authors acknowledge the generous support from the National Science Foundation in the United States. AKYL would like to gratefully acknowledge the sup- port from the National Science Foundation through the Grants NSF PHY-1912594 and NSF PHY-2207758. JCLC acknowledges support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. HF gratefully acknowledges the support of the Canadian Institute for Theoretical Astrophysics (CITA) National Fellowship and the Japan Society for the Promotion of Science (JSPS) Postdoctoral Fellowships for Research in Japan. JME is supported by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 INTERACTIONS, and by VILLUM FONDEN (grant no. 53101 and 37766).
The authors are also grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org) (Vallisneri et al., 2015), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-0757058. HF is supported by a CITA National Fellowship. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes. This paper carries LIGO Document Number LIGO-P1900059.